Project

General

Profile

NDOS Commissioning Meeting 2011/07/18

Attendance: Jon, Susan, Chad, Martin, Raphael, Minerba, Tom, Carl B, Zeb, Jan, Nick, Mat, Andrew, Denis, Gary, Rick, Greg, Athans, Peter, Athans

Agenda:

  1. Beam & Detector Update
  2. DAQ Status & Plans
  3. DCS Status & Plans
  4. AOB

Beam & NDOS Update

  • Jon spoke with Sam Childress re: NT02 preparation last week. Everything seems to be going according to schedule, so we may get beam back by the end of the week.
  • Booster is down due to heat problem in MI-12, could be down for a while (day to days)
  • Detector ran fine over the weekend.
  • Mat:
    • developing a fixture to attach vacuum to optical connection; is able to pull 2 psi, but something in the system was leaking on Friday so he's looking into that.
    • some results from balloon tests, but currently inconclusive. Looks like balloon seals are inconsistent.
    • should have revised APD installation by end of today
  • Denis & Andrew modified DSO scan to run continuously and compute various metrics.that is basically an oscilloscope.
  • Gary: is the HV-noise problem correlated to surface conditions; one should look at individual pixels
  • Jon: let's try to determine rise time
  • Gary & Andrew: more specifically, we'll do an extended scan and watch RMS (eg) drop over time.

DAQ Status & Plans

  • Timing (Andrew): we looked to be timed in at the end of last week, was planning to check this morning but there are no accelerator signals.
  • Neil and Greg are working on adding capability to decode other T-clock signals so we're not strapped to using just beam-based signals. Should be ready today.
  • Will need to take TDUs offline today to reflash them.
  • Andrew: uncovered bug on late Friday; when a run automatically roles over, the all-stream for the new run is empty. Shifters should stop and start a run (nicely).

DCS Status & Plans

  • Athans:
    • not much since last week
    • please remind shifters to give feedback of DCS system to Athans

AOB

  • Jon: will be out of the country Wed-Sat. He'll be in email contact but out of phone contact. Peter S. will available to assist with day-to-day planning, Mark Messier will run the Commissioning meeting on Thursday.
  • Gary: Denis has been looking at electronic noise. Any report? Denis: last week tried to look at those FEBs that still have TECC noise and try to swap TECCs from bad to good and vice-versa to see if the noise is somehow generated in TECCs. Results from 9 tests, 4 did switch from bad to good and good to bad. Denis will try another swap. Gary: is it the same set of pixels that are bad? Denis: not sure, didn't track.